OpenStudy (anonymous):

True or False? If the sides of a square are decreased by 50%, then the area decreases by 75%.

OpenStudy (anonymous):

Area of square=s^2 where s is side of square side is decreased by 50% =0.5S substitute 0.5s in place of s in area formula then u will get answer

OpenStudy (anonymous):

s=.25

OpenStudy (anonymous):

Area=(0.5s)^2=0.25s^2 before the area=s^2 s^2 is decreased to 0.25s^2

OpenStudy (anonymous):

So area is not decreased by 75% but 25%?

OpenStudy (anonymous):

oh wait it would be 50% if squared so it would be true...